OpenStudy (inkyvoyd):

Would this system of differential equations accurately model earth's trajectory around the sun, given that the sun's mass is M, the gravitational constant is g, and the sun is modeled to be stationary at origin (0,0)?

OpenStudy (inkyvoyd):

\(\Huge \frac{d^2x}{dt^2}=\frac{-gMx}{x^2+y^2}\) \(\Huge \frac{d^2y}{dt^2}=\frac{-gMy}{x^2+y^2}\)

OpenStudy (inkyvoyd):

with initial conditions x(0)=1.49*10^11 (distance from earth to sun), y(0)=0 y'(0)=30000 (speed of earth's orbit on average in m/s) x'(0)=0 gM=(gravitational constant g muliplied by mass of the sun)=6.67e-11*1.989e30=1.327e20

OpenStudy (fibonaccichick666):

wouldn't the second deriv model the acceleration?

OpenStudy (inkyvoyd):

yeah, taking F=gmM/r^2, I got the acceleration of earth, a=F/m or gM/r^2 since r is just sqrt(x^2+y^2) I got r^2=(x^2+y^2), and I took component forms of x and y for each of the equations

OpenStudy (inkyvoyd):

supposedly if I solve for x(t) and y(t) numerically or analytically I could get some sort of trajectory, correct?

OpenStudy (fibonaccichick666):

I would just attempt to integrate wrt t and come up with the vector that way... I think

OpenStudy (inkyvoyd):

well, solving the system is not too much an issue for me, but I need to make sure that I've come up with something without any major mistakes

OpenStudy (fibonaccichick666):

well F=Ma to derive a seems fine the substitution for gM/r^2 not sure, you seem to be missing the mass of the earth which would modify the force if I recall correctly. The x and y seems applicable here though

OpenStudy (inkyvoyd):

I thought the mass of the earth cancelled out with the force because ma=Mmg/r^2 so a=Mg/r^2. Does that look right to you?

OpenStudy (fibonaccichick666):

I'm rusty on physics, I need the eq for the force, and isn't that the law for gravitation? wouldn't that be the acceleration?

OpenStudy (inkyvoyd):

oh, I think newton's law of gravitation is that F=gmM/r^2, and newton's second law is F=ma. so I combined the two to get a (of earth)=F/m=gM/r^2, there's also an acceleration for the sun but I think it's negligible.

OpenStudy (fibonaccichick666):

btw use capital G for the constant. Now are you certain on that law of gravitation?

OpenStudy (fibonaccichick666):

yea i just checked it too

OpenStudy (fibonaccichick666):

hmm

OpenStudy (inkyvoyd):

I feel like that x in the numerator doesn't belong there, but I don't know what I should do with it.

OpenStudy (fibonaccichick666):

k so ma=GmM/r^2 a=GM/r^2 r^2=GM/a x^2+y^2= some number implies a perfect circle no?

OpenStudy (fibonaccichick666):

well, then the path is the eq listed above. isn't it?

OpenStudy (inkyvoyd):

Wait, I'm mistaken entirely. My assumption is more so based upon the numbers that I'm using than for the system of equations here. I mean, I'm assuming the average velocity of the earth is around 30km/s, and that the distance the earth is from the sun is some constant, but really I'm trying to reproduce the orbital results (I should see some ellipses)

OpenStudy (fibonaccichick666):

right, but using your method we get circles

OpenStudy (inkyvoyd):

yeah - I intend to take data like the distance 147,098,291 km as the perihelion of the orbit, and 30,300m/s as the speed at perihelion

OpenStudy (inkyvoyd):

since the earth's orbit is elliptical I'm hoping I can see some sort of overlapping ellipses after solving the system.

OpenStudy (inkyvoyd):

I think I'll play around with this a bit more and see if I get anything conclusive. Thanks for your time :)

OpenStudy (inkyvoyd):

got it, my final input was NDSolve[{x''[t] == -grav*x[t]/(((x[t])^2 + (y[t])^2)^(3/2)), y''[t] == -grav*y[t]/(((x[t])^2 + (y[t])^2)^(3/2)), x[0] == 147098291000, y[0] == 0, x'[0] == 0, y'[0] == 30300 }, {x, y}, {t, 100000000}] with grav = 1.327*10^20

OpenStudy (fibonaccichick666):

nice! good luck, sorry i wasn't much help

OpenStudy (inkyvoyd):

tuns out the system I was supposed to be using was \(\Huge \frac{d^2x}{dt^2}=\frac{-gMx}{(x^2+y^2)^{\frac{3}{2}}}\) \(\Huge \frac{d^2y}{dt^2}=\frac{-gMy}{(x^2+y^2)^\frac{3}{2}}\)

OpenStudy (fibonaccichick666):

yucky